OpenStudy (anonymous):

Find all solutions to the equation 2sin^2x-sinx=1

OpenStudy (solomonzelman):

Say, let Sinx=a Solve for a, and then substitute what you get for a, into a=Sinx

OpenStudy (solomonzelman):

\[2\sin^2x-sinx=1~~~~~~->~~~~~2a^2-a=1\]\[2a^2-a-1=0\] can you solve this quadratic equation?